Find the acceleration and the centripetal force of a car that rounds the curved of radius of 50 meters having a speed of 8 km/h.
1
Expert's answer
2020-11-19T08:27:47-0500

The acceleration of the circular motion is given by:


"a = \dfrac{v^2}{R}"

where "v = 8km/h = \dfrac{40}{13}m/s" is the speed and "R = 50m" is the radius. Thus, obtain:


"a = \dfrac{v^2}{R}=\dfrac{40^2}{16^2\cdot 50} = 0.125 m/s^2"

Answer. 0.125 m/s^2